Systems Staff Meeting Notes 05/06/02

Systems Staff Meeting Notes
May 6, 2002

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Harris, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. Cynthia got tickets for the new Star Wars movie to be released on May 16.
  2. Bill read where the IT community is projected to lose $300 million nationwide due to work outages – people taking time off to see the movie.
  3. Geoffrey went to try to see Spiderman on a lark Saturday evening and found that it was sold out until Tuesday.
  4. Tomorrow is the Kickoff & Systems Analysis of ALEPH – Ex-Libris. Twenty-six people will be in attendance – from Smathers, Law and the Health Science Center. Rich will be in attendance the whole time. Bill will at least be attending the initial session and will see how it goes from there. The meetings will be dealing with how to configure and migrate the current NOTIS system into AELPH. Implementation is slated for the next fiscal year. The original version of the system has not even been released yet. Our load (as well as other universities) will increase. Also, look for consulting inquiries. Expect complete angst from staff – NOTIS is the only automated system anyone has used for the last twenty years. The Dean’s Menu should be converted to the Web in the future: we may be able to get rid of 3270 emulation in two years time.
  5. SEMCO still in our closets? Unknown. They’ve received word about locking up when done. If anyone notices the consoles in Facilities unlocked and unattended, lock them up.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Order for 30 more workstations upcoming – no allocation list yet.
  2. Secondary server for MSL (backup for Web, SQL, mail) to be installed.
  3. Trend anti-virus for server to be ordered.
  4. Wiring order is out for the MSL project.
  5. Optical mice order (80) is not here yet.
  6. UPS order is out – 15 big ones.
  1. Diana Hagan has accepted a position as the IT coordinator at IFAS. This will cause disruption within Resource Services due to the internal hiring freeze.
  2. The first librarian liaison will be Steve Carrico.
  1. McAfee is off NT8.
  2. Statistics server went down again on Friday. Will restarted it.
  3. Kleez virus is driving us nuts. It is coming in at double the SirCam rate.
  4. China Syndrome – attack threats originating in China are supposed to happen the first week of May. Friday one of the library staff received a one-word email: “test”. This bloomed to over 15 more over the weekend. Could be a test DOS attack, perhaps some servers have been taken over. There is a known vulnerability with the IP router that could crash the Internet. Please notify Bill if you hear of anyone receiving multiple messages from any one source. He is collecting the info to send to NERDC.
  5. Dell server installation scheduled for this afternoon in the new server room (tape vault). NT2R will need to be rebooted a number of times during the process. Should happen after 4PM. Notice sent out to All Staff last Thursday.
  6. Secure System (lock software) in Facilities will be upgraded to W2K tomorrow.
  7. Master browse – Microsoft told Will to turn on File & Print on all workstations that we want to show up in the Network Neighborhood. Will wants to wait until the PDC is converted to Active Directory to see if that improves the situation.
Software Projects Status
  1. Current Net – ready for switchover. Back and forward all work. Has been re-demoed to Bobbie. It works pretty well. Winston needs one day with no updating to current net in order to switch it over. Will start at 1PM today. Bobbie requests notification when it’s ready to go.
  2. ILLIAD – ILL people are very busy right now – full service ILLIAD starts today.
  3. New project: SysRequest – putting the hardware & software request process online. It is currently in the development stage – Cynthia is working on screens.
  4. Working on mods for several items: Bryant, Newspaper Collection, Florida History Maps, File Sort and DLC.
DP Coordinator
  1. No students this week.
  2. Three brand new students start next Monday: Lori Kress, Nicole Holt and Latoya Daniels.
  3. Fire suppression system inspection in tape vault is scheduled for today.

Back to the Systems Homepage

Last updated May 14, 2002
by Debra Harris